OpenStudy (anonymous):

real or imaginary numbers 8x^3=1

OpenStudy (anonymous):

we have : 8x=1 x=1/8 Got it? Now you say : 1/8 is reall or not ?;)

OpenStudy (anonymous):

@Tim7884

OpenStudy (anonymous):

no

OpenStudy (anonymous):

look at it again E x^3 = 1/8

OpenStudy (anonymous):

tim you want help?

OpenStudy (anonymous):

But in fairness you dont need to solve for x to determine if it is real or imaginary, but it helps knowing how.

OpenStudy (solomonzelman):

\[8x^3=1\]\[8x=1\]\[x=1/8\]\[----->\]he is right!

OpenStudy (zarkon):

that equation has 3 solutions. One real and two imaginary.

OpenStudy (zarkon):

aslo, \(\dfrac{1}{8}\) is not one of the solutions.